# Enhancing Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems

### v1

@@ Abstract
Machine translation focuses mainly on high-resource languages (HRLs), while low-resource languages (LRLs) like Taiwanese Hokkien are relatively under-explored. This study aims to address this gap by developing a dual translation model between Taiwanese Hokkien and both Traditional Mandarin Chinese and English. We employ a pre-trained LLaMA2-7B model specialized in Traditional Mandarin Chinese to leverage the orthographic similarities between Taiwanese Hokkien Han and Traditional Mandarin Chinese. Our comprehensive experiments involve translation tasks across various writing systems of Taiwanese Hokkien and between Taiwanese Hokkien and other HRLs. We find that the use of a limited monolingual corpus also further improve the model’s Taiwanese Hokkien capabilities. We then utilize our translation model to standardize all Taiwanese Hokkien writing systems into Hokkien Han, resulting in further performance improvements. Additionally, we introduce an evaluation method incorporating back-translation and GPT-4 to ensure reliable translation quality assessment even for LRLs. The study contributes to narrowing the resource gap for Taiwanese Hokkien and empirically investigates the advantages and limitations of pre-training and fine-tuning based on LLaMA 2.
@@
Enhancing Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems
| Bo-Han Lu, Yi-Hsuan Lin, En-Shiun Annie Lee, Richard Tzong-Han Tsai††thanks: Corresponding author: thtsai@g.ncu.edu.tw |
| Department of Computer Science and Information Engineering, National Central University, Taiwan |
| Department of Computer Science, Faculty of Science, University of Toronto |
| Center for GIS, Research Center for Humanities and Social Sciences, Academia Sinica, Taiwan |
@@ 1. Introduction
Machine translation (MT), as a crucial subfield of natural language processing (NLP), serves a vital role in overcoming language barriers by translating more texts into the desired language. However, current MT systems predominantly cater to high-resource languages (HRLs), posing significant challenges for low-resource languages (LRLs). Specifically, Taiwanese Hokkien, which is mainly spoken in Taiwan, southern China and a number of countries in Southeast Asia (Ding, 2016a), faces unique issues owing to historical factors (Ding, 2016b) and a persistent absence of standardized writing systems. These factors lead to an extra layer of complexity by introducing inconsistent corpus hinders the development of NLP research and data-hungry translation models for this language.
@@
With the recent advancement of large language models (LLMs) like BLOOM (Scao et al., 2022), ChatGPT and LLaMA (Touvron et al., 2023a), these models have demonstrated their capabilities across various multilingual NLP tasks, including translation tasks (Jiao et al., 2023; García et al., 2023; Yang et al., 2023; Xu et al., 2023). Despite these advancements, state-of-the-art LLMs leave room for improvement in translation tasks, particularly for languages that are considerably removed from HRL (Jiao et al., 2023; Hendy et al., 2023). This study employs a pre-trained LLaMA 2 (Touvron et al., 2023b) model specialized in Mandarin Chinese (ZH), aiming to leverage the orthographic similarities between HAN and ZH to develop a translation model capable of translating between different writing systems of Taiwanese Hokkien as well as between Taiwanese Hokkien and other HRLs like ZH and English.
@@
| English | (EN) | Latin | These few days the weather has not been good. |
@@
For reliable automatic evaluation of translation results, in addition to BLEU score (Papineni et al., 2002) and chrF++ (Popović, 2017) metrics, we modify Kocmi and Federmann (2023)’s evaluation prompt and incorporates a back-translation method so that GPT-4 (OpenAI, 2023) can make reliable evaluations even if the target language is a LRL.
We plan to release a translation system that includes HAN, POJ, ZH and English. We anticipate that this will serve as a reliable translation tool for the public community, and foster the generation of diverse datasets for Taiwanese Hokkien.
@@
Develop and release the first dual translation model for Taiwanese Hokkien, thereby narrowing the resource gap for this low-resource language.
@@ 2. Background
Taiwanese Hokkien, also referred to Hokkien, Hoklo, Taigi, Southern Min, or Min Nan, is a unique subset of the Southern Min dialects. Sharing a common linguistic heritage with the Fujian dialects, Taiwanese Hokkien has undergone an independent evolution influenced by a number of external factors, including indigenous languages and the colonial legacies of the Dutch and Japanese (Liao et al., 2020). In the following discussion, this dialect will be referred to as Hokkien for the purpose of simplicity.
Although Hokkien ranks as the second most common language spoken in Taiwan, recent census222https://www.stat.gov.tw/public/Data/
@@ 2.1. Writing System Diversity in Hokkien
The writing systems of Hokkien can be divided into three main groups. The first is Hokkien Han (HAN), which is based on Traditional Chinese with additional characters, followed by Latin script systems such as Tâi-lô (TL) and Pe̍h-ōe-jī (POJ). In addition, a hybrid system known as Hàn-lô (HL) combines elements of both systems. Since 2009, an official orthography for HAN has been established and is currently used in educational systems. However, due to its relatively recent standardization, corpus resources of HAN are scarce compared to other systems. On the other hand, POJ, which was introduced by missionaries in the 19th century, has a substantial amount of digitized historical texts and thus provides a rich corpus of Hokkien writings. Moreover, TL, an adaptation of POJ, maintains a systematic correspondence with it. HL, with its mixture of Latin script and Chinese characters, exhibits considerable variance across textual resources due to the lack of a uniform standard for determining the use of Latin script or Chinese characters. Therefore, in this study, HL corpus is considered for training purposes, but is excluded from translation evaluations.
@@ 2.2. Semantic Divergence of Shared Chinese Characters in HAN and ZH
Despite the commonality of Chinese characters between HAN and ZH, many homographs differ semantically. For example, the term “手指” translates to “finger” in ZH, while it means “ring” in HAN. Moreover, common HAN terms often correspond to rarely used Chinese characters in ZH, such as ‘覕’ (hide) and ‘啉’ (drink). As a result, training a reliable translation model capable of translating between these two languages under low-resource conditions still remains challenge.
@@ 3.1. Large Language Models in Translation
LLMs have recently made remarkable progress in translation tasks due to their robust language understanding capabilities from pre-training on massive corpora. In the field of applying LLMs to translation tasks, Moslem et al. (2023); Lin et al. (2022); Zhu et al. (2023); Zhang et al. (2023a); Vilar et al. (2023); García et al. (2023) attempted to use in-context learning (ICL) (Brown et al., 2020) to enhance the translation capabilities of LLMs. Their study demonstrated how the pattern of in-context learning, the selection of few-shot sentences, and their quantity could impact the translation results. Zhang et al. (2023b); Yang et al. (2023); Li et al. (2023) tried to enhance the translation abilities of LLMs through instruction-tuning (Ouyang et al., 2022) with small amounts of parallel data. Li et al. (2023) demonstrated 3 BLEU score on average advancements of multilingual translation when compared to the ICL method. Some research (Hendy et al., 2023; Jiao et al., 2023) has indicated that LLMs may have limited translation abilities because their language skills are largely shaped by training in English-centered texts. The translation proficiency of LLMs is often significantly limited when translating languages that are not linguistically close to English (Li et al., 2023). As a result, Yang et al. (2023) and Li et al. (2023) have included different translation languages for monolingual training. Xu et al. (2023)’s latest findings also adopt monolingual training before fine-tuning translation task. Using medium-sized models with 7B and 13B parameters, they surpassed GPT-3.5 and NLLB-54B (Team et al., 2022) in various translation tasks.
@@ 3.2. Neural Machine Translation in Hokkien
Due to the scarcity of training data, neural machine translation (NMT) for low-resource languages such as Hokkien faces unique challenges. Liao et al. (2022) has compiled a dataset for Hokkien speech recognition, which not only contributes to speech-related research, but also benefits NMT through the transcribed parallel data. The techniques of transfer learning and cross-lingual models offer possible ways to improve Hokkien NMT systems. Lu et al. (2022) investigates translation task between ZH, Hokkien code-mixing language and ZH. They apply transfer learning to utilize the knowledge pre-trained on ZH by XLM (Conneau and Lample, 2019), and develop a method to synthesize a code-mixing translation parallel dataset to achieve better translation results between code-mixing language and ZH. To the best of our knowledge, we are the first to explore the application of large language models to dual translation for Hokkien while considering both Latin script and Chinese character writing systems of Hokkien.
@@ 4.1.1. Monolingual Datasets
Our dataset primarily comprises web articles in Hokkien collected from diverse internet sources including scripts from Hokkien recitation contest333http://ip194097.ntcu.edu.tw/longthok/longthok.asp, lyrics of Hokkien songs shared in Facebook communities444https://www.facebook.com/groups/922800454445724, and web-scraped articles covering various domains. The dataset also incorporates religious articles in Hokkien, content from Wikipedia, and Hokkien elementary school textbooks.
Additionally, we include subtitles from Hokkien television programs. Given that these subtitles often lack paragraph breaks and punctuation, we employed GPT-3.5-turbo to refine the textual structure, rendering it more akin to standard articles.
@@ 4.1.2. Parallel Datasets
For fine-tuning data, we collected the datasets from example sentences in the Hokkien dictionary555https://sutian.moe.edu.tw/zh-hant/, religious texts666http://pojbh.lib.ntnu.edu.tw/script/index.php, and technical terms777https://stti.moe.edu.tw/?lang=sutgi. For those texts that have corresponding parallel sentences in ZH, we utilized GPT-3.5-turbo to translate them into English. Further details regarding the writing system and other fundamental features for the datasets are provided in Table 3.
@@ 4.2. Model Training
To leverage the shared Chinese character system between ZH and HAN, we chose TAIDE-7B888https://taide.tw/ as our base model. Enriched with an additional 24k Chinese character tokens and pre-trained on a 1.7B-token Traditional Chinese corpus, TAIDE-7B serves as an extension of LLaMA 2 (Touvron et al., 2023a), enhanced comprehension of Taiwan-specific Traditional Chinese terms.
@@
Although the TAIDE-7B model contains a large amount of Chinese characters, it still lacks coverage for Hokkien Latin scripts and some rarely used Chinese characters specific to Hokkien. To address this, we further extend the vocabulary by training a sentence-piece (Kudo and Richardson, 2018) tokenizer on monolingual Hokkien corpora and merge it back to the original one. Specifically, the vocabulary was extended with an additional 130 Chinese character tokens for HAN and 1876 Latin script tokens for POJ, resulting in a final vocabulary size of 58,505.
@@
We conducted continued pre-training on monolingual Hokkien corpora across all writing systems. We followed procedures from Chinese-LLaMA-2 Cui et al. (2023) using Low-Rank Adaptation (LoRA) Hu et al. (2021) with gradient checkpoint to reduce computational cost, and trained 18 epochs to avoid undesirable outputs. Given the rule-based transformation between the POJ and TL systems, we converted all TL in the corpus to POJ using an existing tool999https://github.com/i3thuan5/KeSi, streamlining the model training across these two Latin script writing systems.
@@
The label [TRANS] denotes translation, X and Y are the source and target sentences, respectively. The label [target_lang] indicates the target language for the translation. During fine-tuning, each language pair in the parallel data was fixed to 17,872 instances, including HAN-ZH, HAN-EN, POJ-ZH, POJ-EN, and POJ-HL, and each model was trained for one epoch.
@@ 4.3.1. Two Translation Testing Datasets
We used iCorpus101010https://github.com/Taiwanese-Corpus/icorpus_ka1_han3-ji7 as our testing data. It comprises news articles from various domains and includes both HAN and POJ. However under human evaluation, the HAN section of iCorpus contains a considerable number of lexical inaccuracies. To cover this issue in the test set, we first selected terms in HAN that deviate significantly from ZH, using their frequency as a selection criterion to ensure the translation difficulty. Based on this criterion, we sampled the top 100 sentences where HAN appears frequently and manually corrected their lexical according to official orthography. This resulted in a subset we named iCorpus-100. Due to the size limitation of iCorpus-100, we integrated an additional data from TAT (Liao et al., 2020), specifically to enhance the evaluation of ZHHAN and ENHAN translations. TAT comprises 2,661 parallel sentences sourced from the Taiwanese Across Taiwan speech recognition competitions. Using the same process as the training set, the EN side of the parallel data were generated through translation from ZH using GPT-3.5-turbo. Neither of these test data was used to continued pre-training and fine-tuning.
@@ 4.3.2. Evaluation Metrics
In order to conduct a comprehensive evaluation of our translation models, we used four different metrics: BLEU score (Papineni et al., 2002), chrF++ (Popović, 2017), and two additional GPT-based metrics. Unlike BLEU, chrF++ includes evaluations at the character level. This lead to significant difference between these two metrics when the target language is Latin script, especially POJ. We attribute this to the accent variations in POJ that may alter one or two characters within a word. Such accent-induced variations were not normalized in out corpus. As a result, the model may produce accent inconsistencies with the ground truth, leading to significant penalties in BLEU score, but relatively small drop in chrF++.
Though widely used in translation assessment, BLEU and chrF++ mainly focus on lexicon-level granularity and do not provide a comprehensive evaluation of overall translation quality. In particular, these measures are limited when applied to large language models, which often require more nuanced methods of evaluation (Hendy et al., 2023). Kocmi and Federmann (2023) indicates that GPT-4 holds promise for accurate evaluation of the translation outputs. Therefore, we employ GPT-4 to assess the generated translations, ranging from 0 to 100, by providing both the model’s translation and a reference. However, GPT-4’s ability to comprehend Hokkien is restricted, therefore to overcome this constraint, we implement back-translation for the model’s output when translating from ZH or English into Hokkien. The back-translated output is then compared with the source sentence for evaluation. This method is not suitable for scenarios where both the source and target sentences are in Hokkien. For the evaluation of translation scores from GPT-4, we conducted a human qualitative analysis on 20 samples from each of five score intervals, totaling 100 examples. The qualitative findings based on these analyses are summarized in 4.
After the evaluation with GPT-4, we calculate the GPT-4 score by taking the average of each translation result. Since translations scoring 80 or above closely approximate the original sentence’s meaning, we consider them as correct translations, which are used to compute the translation accuracy, named as GPT-4 accuracy.
@@
80-100 |
The translation matches the source in meaning, using identical terms or appropriate synonyms. The overall grammar is correct, and the meaning flows smoothly. |
60-79 |
Errors occur with the translation of proper nouns (e.g., names) or misuse of nouns or verbs, causing slight shifts in meaning. At times, some information may be omitted in the translation or there may be over-translations. |
40-59 |
Parts of the translation are unrelated to the source text. Misuse of verbs and nouns in inappropriate contexts occurs more frequently. |
20-39 |
Most of the translation has deviations from the source text, though occasionally it still conveys the basic idea. Grammar and sentence structures are often incorrect, making the translated content challenging to understand. |
0-19 |
The translation is almost entirely unrelated to the source text. Sentences are not fluent, with frequent major grammatical and structural errors. |

### v2

# Enhancing Taiwanese Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems

@@ Abstract
Machine translation focuses mainly on high-resource languages (HRLs), while low-resource languages (LRLs) like Taiwanese Hokkien are relatively under-explored. The study aims to address this gap by developing a dual translation model between Taiwanese Hokkien and both Traditional Mandarin Chinese and English. We employ a pre-trained LLaMA 2-7B model specialized in Traditional Mandarin Chinese to leverage the orthographic similarities between Taiwanese Hokkien Han and Traditional Mandarin Chinese. Our comprehensive experiments involve translation tasks across various writing systems of Taiwanese Hokkien as well as between Taiwanese Hokkien and other HRLs. We find that the use of a limited monolingual corpus still further improves the model’s Taiwanese Hokkien capabilities. We then utilize our translation model to standardize all Taiwanese Hokkien writing systems into Hokkien Han, resulting in further performance improvements. Additionally, we introduce an evaluation method incorporating back-translation and GPT-4 to ensure reliable translation quality assessment even for LRLs. The study contributes to narrowing the resource gap for Taiwanese Hokkien and empirically investigates the advantages and limitations of pre-training and fine-tuning based on LLaMA 2.
@@
Enhancing Taiwanese Hokkien Dual Translation by Exploring
and Standardizing of Four Writing Systems
| Bo-Han Lu†, Yi-Hsuan Lin†, En-Shiun Annie Lee‡§, Richard Tzong-Han Tsai†∗C††thanks: CCorresponding author: thtsai@g.ncu.edu.tw |
| †Department of Computer Science and Information Engineering, National Central University, Taiwan |
| ‡Department of Computer Science, Faculty of Arts and Science, University of Toronto |
| §Computer Science Program, Faculty of Science, Ontario Tech University |
| ∗Center for GIS, Research Center for Humanities and Social Sciences, Academia Sinica, Taiwan |
@@ 1. Introduction
Machine translation (MT), as a crucial subfield of natural language processing (NLP), serves a vital role in overcoming language barriers by translating more texts into the desired language. However, current MT systems predominantly cater to high-resource languages (HRLs), posing significant challenges for low-resource languages (LRLs). Specifically, Taiwanese Hokkien, which is mainly spoken in Taiwan, southern China and a number of countries in Southeast Asia (Ding, 2016a), faces unique issues owing to historical factors (Ding, 2016b) and a persistent absence of standardized writing systems. These factors lead to an extra layer of complexity by introducing inconsistent corpora, which hinders the development of NLP research and data-hungry translation models for this language.
@@
With the recent advancement of large language models (LLMs) like BLOOM (Scao et al., 2022), ChatGPT and LLaMA (Touvron et al., 2023a), these models have demonstrated their capabilities across various multilingual NLP tasks, including translation tasks (Jiao et al., 2023; García et al., 2023; Yang et al., 2023; Xu et al., 2023). Despite these advancements, state-of-the-art LLMs leave room for improvement in translation tasks, particularly for languages that are considerably removed from HRL (Jiao et al., 2023; Hendy et al., 2023).
This study employs a pre-trained LLaMA 2 (Touvron et al., 2023b) model specialized in Mandarin Chinese (ZH), aiming to leverage the orthographic similarities between HAN and ZH to develop a translation model capable of translating between different writing systems of Taiwanese Hokkien as well as between Taiwanese Hokkien and other HRLs like ZH and English.
@@
| English | (EN) | Latin | It’s been bad weather these days. |
@@
For reliable automatic evaluation of translation results, in addition to BLEU score (Papineni et al., 2002) and chrF++ (Popović, 2017) metrics, we modify Kocmi and Federmann (2023)’s evaluation prompt and incorporates a back-translation method (Rapp, 2009) so that GPT-4 (OpenAI, 2023) can make reliable evaluations even if the target language is a LRL.
We plan to release the translation model that includes HAN, POJ, ZH and English. We anticipate that this will serve as a reliable translation tool for the public community, and foster the generation of diverse datasets for Taiwanese Hokkien.
@@
Develop and release the first dual translation model for Taiwanese Hokkien, thereby narrowing the resource gap for this low-resource language.222The model and other related resources are available at https://github.com/lbh0830/TW-Hokkien-LLM.
@@ 2. Background
Taiwanese Hokkien, also known as Hokkien, Hoklo, Taigi, Southern Min, or Min Nan, is a unique subset of the Southern Min dialects. Sharing a common linguistic heritage with the Fujian dialects, Taiwanese Hokkien has undergone an independent evolution influenced by a number of external factors, including indigenous languages and the colonial legacies of the Dutch and Japanese (Liao et al., 2020). In the following discussion, this dialect will be referred to as “Hokkien” for the purpose of simplicity.
Although Hokkien ranks as the second most common language spoken in Taiwan, recent census333https://www.stat.gov.tw/public/Data/
@@ 2.1. Writing System Diversity in Hokkien
[…]
@@ 2.2. Semantic Divergence of Shared Chinese Characters in HAN and ZH
Despite the commonality of Chinese characters between HAN and ZH, many homographs differ semantically. For example, the term “手指” translates to “finger” in ZH, while it means “ring” in HAN. Moreover, common HAN terms often correspond to rarely used Chinese characters in ZH, such as ‘覕’ (hide) and ‘{CJK}UTF8gbsn啉’ (drink). As a result, training a reliable translation model capable of translating between these two languages under low-resource conditions still remains challenges.
@@ 3.1. Large Language Models in Translation
LLMs have recently made remarkable progress in translation tasks due to their robust language understanding capabilities from pre-training on massive corpora. In the field of applying LLMs to translation tasks, Moslem et al. (2023); Lin et al. (2022); Zhu et al. (2023); Zhang et al. (2023a); Vilar et al. (2023); García et al. (2023) attempted to use in-context learning (ICL) (Brown et al., 2020) to enhance the translation capabilities of LLMs. Their study demonstrated how the pattern of in-context learning, the selection of few-shot sentences, and their quantity could impact the translation results. Zhang et al. (2023b); Yang et al. (2023); Li et al. (2023) tried to enhance the translation abilities of LLMs through instruction-tuning (Ouyang et al., 2022) with small amounts of parallel data. Li et al. (2023) demonstrated BLEU score on average advancements of multilingual translation when compared to the ICL method. Some research (Hendy et al., 2023; Jiao et al., 2023) has indicated that LLMs may have limited translation abilities because their language skills are largely shaped by training in English-centered texts. The translation proficiency of LLMs is often significantly limited when translating languages that are not linguistically close to English (Li et al., 2023). As a result, Yang et al. (2023) and Li et al. (2023) have included different translation languages for monolingual training. Xu et al. (2023)’s latest findings also adopt monolingual training before fine-tuning translation tasks. Using medium-sized models with 7B and 13B parameters, they surpassed GPT-3.5 and NLLB-54B (Team et al., 2022) in various translation tasks.
@@ 3.2. Neural Machine Translation in Hokkien
Due to the scarcity of training data, neural machine translation (NMT) for low-resource languages such as Hokkien faces unique challenges. Liao et al. (2022) has compiled a dataset for Hokkien speech recognition, which not only contributes to speech-related research but also benefits NMT through the transcribed parallel data. The techniques of transfer learning and cross-lingual models offer possible ways to improve Hokkien NMT systems. Lu et al. (2022) investigates translation task between ZH, Hokkien code-mixing language and ZH. They apply transfer learning to utilize the knowledge pre-trained on ZH by XLM (Conneau and Lample, 2019), and develop a method to synthesize a code-mixing translation parallel dataset to achieve better translation results between the code-mixing language and ZH. To the best of our knowledge, we are the first to explore the application of large language models to dual translation for Hokkien, accommodating both its Latin script and Chinese character writing systems.
@@ 4.1.1. Monolingual Datasets
Our corpus primarily comprises web articles in Hokkien collected from diverse internet sources including scripts from Hokkien recitation contests444https://han-tsi5.knsh.com.tw/Resource.asp?T=MM, lyrics of Hokkien songs shared in Facebook communities555https://www.facebook.com/groups/922800454445724, and web-scraped articles covering various domains. The corpus also incorporates religious articles in Hokkien, content from Wikipedia, and Hokkien elementary school textbooks.
Additionally, we include subtitles from Hokkien television programs. Given that these subtitles often lack paragraph breaks and punctuation, we employed GPT-3.5-turbo666The term “GPT-3.5-turbo” in this paper specifically refers to the version “gpt-3.5-turbo-0613”. to refine the textual structure, rendering it more akin to standard articles.
@@ 4.1.2. Parallel Datasets
For fine-tuning data, we collected the datasets from example sentences in the Hokkien dictionary777https://sutian.moe.edu.tw/zh-hant/, religious texts888http://pojbh.lib.ntnu.edu.tw/script/index.php, and technical terms999https://stti.moe.edu.tw/?lang=sutgi. Due to the lack of direct Hokkien-to-English parallel dataset, for those texts that have corresponding parallel sentences in ZH, we used GPT-3.5-turbo to translate them into English. Further details regarding the writing system and other fundamental features for the datasets are provided in Table 3.
@@ 4.2. Model Training
To leverage the shared Chinese character system between ZH and HAN, we chose TAIDE-7B academic research model101010https://taide.tw/ as our base model. Enriched with an additional 24k Chinese character tokens and pre-trained on a 1.7B-token Traditional Chinese corpus, TAIDE-7B serves as an extension of LLaMA 2 (Touvron et al., 2023a), enhanced comprehension of Taiwan-specific Traditional Chinese terms.
@@
Although the vocabulary of TAIDE-7B model contains a large amount of Chinese characters, it still lacks coverage for Hokkien Latin scripts and some rarely used Chinese characters specific to Hokkien. To address this, we further extend the vocabulary by training a sentence-piece (Kudo and Richardson, 2018) tokenizer on monolingual Hokkien corpora and merge it back to the original one. Specifically, the vocabulary was extended with an additional 130 Chinese character tokens for HAN and 1876 Latin script tokens for POJ, resulting in a final vocabulary size of 58,505.
@@
We conducted continued pre-training on monolingual Hokkien corpora across all writing systems. We followed procedures from Chinese-LLaMA-2 Cui et al. (2023) using Low-Rank Adaptation (LoRA) Hu et al. (2021) with gradient checkpoint to reduce computational cost, and trained 18 epochs to avoid undesirable outputs. Given the rule-based transformation between the POJ and TL systems, we converted all TL in the corpus to POJ using an existing tool111111https://github.com/i3thuan5/KeSi, streamlining the model training across these two Latin script writing systems.
@@
The label [TRANS] denotes translation, X and Y are the source and target sentences, respectively. The label [{target_lang}] indicates the target language for the translation. During fine-tuning, each language pair in the parallel data was fixed to 17,872 instances, including HAN-ZH, HAN-EN, POJ-ZH, POJ-EN, and POJ-HL. Each model was trained for one epoch.
@@ 4.3.1. Two Translation Testing Datasets
We used iCorpus121212https://github.com/Taiwanese-Corpus/icorpus_ka1_han3-ji7, a resource originally created by Academia Sinica and subsequently augmented by communities, as our testing data. It comprises news articles from various domains and includes HAN, POJ and ZH. However under human evaluation, the HAN section of iCorpus contains a considerable number of lexical inaccuracies. To cover this issue in the test set, we first selected terms in HAN that deviate significantly from ZH, using their frequency as a selection criterion to ensure the translation difficulty. Based on this criterion, we sampled the top 100 sentences where HAN appears frequently and manually corrected their lexicons according to official orthography. This resulted in a subset that we named iCorpus-100. Due to the size limitation of iCorpus-100, we integrated an additional data from TAT (Liao et al., 2020), specifically to enhance the evaluation of ZHHAN and ENHAN translations. TAT comprises 2,661 parallel sentences sourced from the Taiwanese Across Taiwan speech recognition competitions. Using the same process as the training set, the EN side of the parallel data were generated through translation from ZH using GPT-3.5-turbo. Neither of these test data was used in continued pre-training and fine-tuning stages.
@@ 4.3.2. Evaluation Metrics
In order to conduct a comprehensive evaluation of our translation models, we used four different metrics: BLEU score (Papineni et al., 2002), chrF++ (Popović, 2017), and two additional GPT-based metrics. Unlike BLEU score, chrF++ includes evaluations at the character level. This lead to significant difference between these two metrics when the target language is Latin script, especially POJ. We attribute this to the accent variations in POJ that may alter one or two characters within a word. Such accent-induced variations were not normalized in our corpus. As a result, the model may produce accent inconsistencies with the ground truth, leading to significant penalties in BLEU score, but relatively small drop in chrF++.
Though widely used in translation assessment, BLEU score and chrF++ mainly focus on lexicon-level granularity and do not provide a comprehensive evaluation of overall translation quality. In particular, these measures are limited when applied to large language models, which often require more nuanced methods of evaluation (Hendy et al., 2023). Kocmi and Federmann (2023) indicate that GPT-4 holds promise for accurate evaluation of the translation outputs. Therefore, we employ GPT-4131313The term “GPT-4” in this paper specifically refers to the version “gpt-4-0613”. to assess the generated translations, ranging from 0 to 100, by providing both the model’s translation and a reference. However, GPT-4’s ability to comprehend Hokkien is restricted. To overcome this constraint, we use the approach inspired by Rapp (2009) to implement back-translation for the model’s output when translating from ZH or English into Hokkien. We then compare the back-translated output with the source sentence for evaluation. Note that this approach is not applicable in situations where both the source and target sentences are in Hokkien. For the evaluation of translation scores from GPT-4, we conducted a human qualitative analysis on 20 samples from each of the five score intervals, totaling 100 examples. The qualitative findings based on these analyses are summarized in Table 4. Additionally, the prompt template for GPT-4 evaluations and the examples corresponding to these five score intervals are detailed in Appendix A.
After evaluating with GPT-4, we calculate the GPT-4 score by taking the average of each translation result. Since translations scoring 80 or above closely approximate the original sentence’s meaning, we consider them as correct translations, which are used to compute the translation accuracy, named as GPT-4 accuracy.
@@
| 80-100 | The translation matches the source in meaning, using identical terms or appropriate synonyms. The overall grammar is correct, and the meaning flows smoothly. |
| 60-79 | Errors occur with the translation of proper nouns (e.g., names) or misuse of nouns or verbs, causing slight shifts in meaning. At times, some information may be omitted in the translation or there may be over-translations. |
| 40-59 | Parts of the translation are unrelated to the source text. Misuse of verbs and nouns in inappropriate contexts occurs more frequently. |
| 20-39 | Most of the translation has deviations from the source text, though occasionally it still conveys the basic idea. Grammar and sentence structures are often incorrect, making the translated content challenging to understand. |
| 0-19 | The translation is almost entirely unrelated to the source text. Sentences are not fluent, with frequent major grammatical and structural errors. |